OpenStudy (anonymous):

7(f-1)=45 help plzz.

OpenStudy (saifoo.khan):

7f - 7 = 45 7f = 45 - 7 7f = 38 f = 38/7

jimthompson5910 (jim_thompson5910):

7(f-1)=45 7f-7=45 7f=45+7 7f=52 f=52/7

OpenStudy (anonymous):

Jim_thompson's answer is correct